solve for x 4=256^3x

Answer by Alan3354(69443)

solve for x 4=256^3x
------------------------
If it's 4=256^(3x)
4 = (4^4)^(3x) = 4^(12x)
12x = 1
x = 1/12
-----------------
If it's 4=256^(3)x
4 = (4^4)^3*x
4 = 4^12 *x
1 = 4^11*x
x = 1/(4^11) =~ 0.0000002.3841858
=======================================
Parentheses are free. Use some.